classify these reactions
2HBr(aq)+Ba(OH){2}(aq)= 2H{2}0(l)+BaBr{2}(aq)

CH{4}(g)+2o{2}(g)=CO{2}(g)+2H{2}O(l)

Cu(s)+FeCl{2}(aq)=Fe(s)+CaCl{2}(aq)

Na{2}S(aq)+2AgNO{3}(aq)+2NaNO{3}(aq)+Ag{2}S(s)
1
Expert's answer
2015-11-11T11:21:33-0500
CH4(g)+2O2(g)=CO2(g)+2H2O(l)

It is burning reaction (oxidations by oxygen).
This reaction is heterogeneous (one of substances is liquid, the others - gases).


2HBr(aq)+Ba(OH)2(aq)= 2H2O(l)+BaBr2(aq)

It is reaction of an exchange (two difficult substances exchange the components).
This reaction is homogeneous (all substances are in water solution).


Cu(s)+FeCl2(aq)=Fe(s)+CaCl2(aq)

It is replacement reaction (iron forces out copper from its salt).
This reaction is oxidation-reduction since copper and iron change oxidation level: iron is oxidized, and copper is restored.
This reaction is heterogeneous (one of substances is firm and not soluble in water, other substances - soluble in water).


Na2S(aq)+2AgNO3(aq) + 2NaNO3(aq)+Ag2S(s)

It is reaction of an exchange (two difficult substances exchange the components).
This reaction is heterogeneous (one of substances is firm and not soluble in water, other substances - soluble in water).